# neurips_2025.bib

### v1

@@ Abstract
Many architectures for neural PDE surrogates have been proposed in recent years, largely based on neural networks or operator learning. In this work, we derive and propose a new architecture, the Neural Functional, which learns function to scalar mappings. Its implementation leverages insights from operator learning and neural fields, and we show the ability of neural functionals to implicitly learn functional derivatives. For the first time, this allows for an extension of Hamiltonian mechanics to neural PDE surrogates by learning the Hamiltonian functional and optimizing its functional derivatives. We demonstrate that the Hamiltonian Neural Functional can be an effective surrogate model through improved stability and conserving energy-like quantities on 1D and 2D PDEs. Beyond PDEs, functionals are prevalent in physics; functional approximation and learning with its gradients may find other uses, such as in molecular dynamics or design optimization.
@@ 1 Introduction
Despite the long history of Hamiltonian mechanics, most current numerical and neural PDE solvers operate within a Newtonian framework. This has been very successful; forces and masses are easy to intuitively understand, and both numerical and neural solvers have found important, practical uses. However, motivated by natural conservation laws and symmetries that arise from Hamiltonian mechanics, we seek to investigate neural PDE surrogates that operate in this framework. We find that the Hamiltonian framework brings inductive biases that improve model performance, especially when conserving energy-like quantities.
Extending Hamiltonian mechanics to neural PDE surrogates requires both theoretical and implementation insights. In particular, most PDEs are used to describe continuum systems (fluids, waves, elastic bodies, etc.); as such, infinite-dimensional Hamiltonian mechanics are needed. Within this setup, the Hamiltonian is defined as a functional that maps from input fields to a scalar, often interpreted as the energy of the system. The variational or functional derivative of the Hamiltonian then contains dynamical information that evolves the system. This introduces new modeling challenges; in particular, no current neural architecture can theoretically model this infinite-dimensional to scalar mapping, let alone have meaningful functional derivatives.
In this paper, we introduce the Neural Functional, an architecture that can provably represent functionals in the linear case and implicitly learn functional derivatives. Through automatic differentiation, its functional derivative can obtained and optimized in a Hamiltonian framework to predict future PDE states. We evaluate the Hamiltonian Neural Functional (HNF) on three PDE systems and find both high accuracy and generalization capabilities, which we hypothesize arise from inductive biases to conserve energy. Beyond predicting PDEs, functionals arise in many physical systems, opening interesting opportunities to apply and improve neural functionals for other meaningful uses. Code and datasets for this work are released at https://github.com/anthonyzhou-1/hamiltonian_pdes.
@@ 2 Background
Many PDEs describe continuum systems, such as waves, fluids, and elastic bodies, where the positions and momenta of discrete bodies are not well defined. Despite this change, continuum systems can still conserve energy and be viewed in a Hamiltonian framework. Firstly, finite-dimensional vectors $\mathbf{q},\mathbf{p}\in\mathbb{R}^{n}$ become generalized to functions $u(x,t)$ defined on continuously varying coordinates $x\in\Omega$ and at a time $t$. The function $u\in\mathcal{F}(\Omega)$ describes the state of the system by assigning a scalar or vector for each coordinate (such as velocity, pressure, etc.). Secondly, the Hamiltonian is generalized from a function $\{\mathcal{H}(\mathbf{q},\mathbf{p}):\mathbb{R}^{2n}\rightarrow\mathbb{R}\}$ to a functional $\{\mathcal{H}[u]:\mathcal{F}(\Omega)\rightarrow\mathbb{R}\}$ that returns a scalar given an input function. Lastly, the symplectic matrix $J$ generalizes to a linear operator $\mathcal{J}$, and the gradient $\nabla_{\mathbf{u}}$ becomes the variational or functional derivative $\frac{\delta\mathcal{H}}{\delta u}\in\mathcal{F}(\Omega)$. This leads to Hamilton’s equations for infinite-dimensional systems, where $\epsilon\in\mathcal{F}(\Omega)$ is an arbitrary test function:
@@
[…]
3 Neural Functionals
@@
The development and underlying theory of neural functionals is a result of the Riesz representation theorem, a fundamental result in functional analysis [kreyszig1991introductory]. We give an informal derivation of neural functionals to provide intuition for the architecture; a full proof is in Appendix C. We start by observing that a functional can be equivalent to the inner product over suitable functions:
@@
Immediately, we can see that the Hamiltonian $\mathcal{H}[u]$ can potentially be expressed as an inner product $\left\langle u,\kappa_{\theta}\right\rangle$ over the input function $u$ and a learnable function $\kappa_{\theta}$. This changes the problem of approximating a functional to the problem of approximating a function, which neural networks can readily achieve through universal approximation theorems [Pinkus_1999, LESHNO1993861].
A common inner product for arbitrary functions $u,v$ in a Hilbert space $H$ involves an integral: $\left\langle u,v\right\rangle=\int_{\Omega}u(x)v(x)\text{d}x$. One can check that this satisfies the basic properties of inner products and is also valid for vector-valued functions $u,v$. Based on this formulation, we propose to parameterize Neural Functionals by the Integral Kernel Functional (IKF):
@@
In this form, we observe a connection between the integral kernel functional and the integral kernel operator that neural operators are built on [kovachki_neuraloperator]. Specifically, the integral kernel operator is given by:
@@
[…]
with quadrature weights $\mu_{i}$. Interestingly, for neural functionals the full Riemann sum can be computed since the sum is only calculated once; in neural operator literature, approximating the integral with a full Riemann sum is usually too costly and the sum is truncated [li2020neuraloperatorgraphkernel, multipole, li2023geometryinformedneuraloperatorlargescale] or represented in Fourier space through the convolution theorem [li2021fourierneuraloperatorparametric]. In our implementation, we evaluate different quadratures but find that the trapezoidal rule is enough to accurately estimate the integral.
To parameterize the kernel, we take the perspective from the Riesz representation theorem. In the linear case, the kernel is simply a function $\kappa_{\theta}(x)\in H$ that maps input coordinates to values in the codomain. This is exactly what a neural field is; therefore, we can inherit the substantial prior work on neural fields [xie2022neuralfieldsvisualcomputing] to effectively parameterize $\kappa_{\theta}$.
One useful concept is the conditional neural field. In particular, the nonlinear kernel $\kappa_{\theta}(x,u(x))$ can be seen as a conditional neural field whereby the kernel output changes based on the input function $u(x)$. This concept can also be further extended to distinguish between local and global conditioning [peng2020convolutionaloccupancynetworks]; in neural field literature, the field can be conditioned on local information $u_{i}=u(x_{i})$ or global information $\mathbf{u}=[u(x_{0}),\ldots,u(x_{n})]$. Conditioning mechanisms can also be implemented in various ways. In the simplest case, the conditional information $u_{i}$ or $\mathbf{u}$ is concatenated to the input coordinate $x$ [mehta2021modulatedperiodicactivationsgeneralizable, park2019deepsdflearningcontinuoussigned, mescheder2019occupancynetworkslearning3d], however, we adopt a approach based on Feature-wise Linear Modulation (FiLM) [perez2017filmvisualreasoninggeneral].
Beyond conditioning mechanisms, an important consideration is the architecture of the kernel itself. A powerful neural field architecture is the sinusoidal representation network or SIREN [sitzmann2020implicitneuralrepresentationsperiodic]. SIRENs have shown good performance not only in fitting a neural field $\kappa_{\theta}$, but also in accurately representing gradients $\nabla\kappa_{\theta}$, which can aid in fitting the functional derivative $\delta\mathcal{H}/\delta u$. With these architectural choices, a layer of the kernel $\kappa^{(l)}_{\theta}$ with local conditioning can be written as:
@@
Hamiltonian Neural Functional
So far, the developed architecture is largely agnostic to the downstream task, only requiring that the input function $u$ be discretized at a set of points $(x_{i},u_{i})$. To use neural functionals as a PDE surrogate model, we introduce the Hamiltonian Neural Functional (HNF). The time evolution of PDEs can be described by Hamilton’s equations, which depend on the operator $\mathcal{J}$ and the functional derivative $\delta\mathcal{H}/\delta u$. In Hamiltonian PDE systems, $\mathcal{J}$ is analytically derived or can be looked up, is guaranteed to be linear, and is usually simple. A common example in 1D is $\mathcal{J}=\partial_{x}$. Therefore, we choose to approximate $\mathcal{J}$ with a finite difference scheme, although it can also be learned with a neural operator. Lastly, we find that $\delta\mathcal{H}/\delta u$ can be computed directly with automatic differentiation. This is empirically shown in Section 4.1, but one way to see this is by considering the linear IKF. The gradient $\nabla_{\mathbf{u}}\sum_{i=1}^{n}\kappa_{\theta}(x_{i})u_{i}\>\mu_{i}\Delta x=[\kappa_{\theta}(x_{1}),\ldots,\kappa_{\theta}(x_{n})]$, up to a constant factor, which is the discretization of the learned function $\kappa_{\theta}(x)$. IN this case, the functional derivative is approximated by an arbitrary function, which is the desired behavior.
Algorithm 1 Training a HNF
@@
2:  $\mathcal{H}_{\theta}\leftarrow\sum_{i=1}^{n}\kappa_{\theta}(x_{i},u_{i})u_{i}\>\mu_{i}\Delta x$
3:  $\frac{\delta\mathcal{H}_{\theta}}{\delta u}\leftarrow\text{autograd}(\mathcal{H}_{\theta},\mathbf{u})$
4:  $\mathcal{L}=||\frac{\delta\mathcal{H}_{\theta}}{\delta u}-\frac{\delta H}{\delta u}||^{2}\>\text{or}\>||\mathcal{J}(\frac{\delta\mathcal{H}_{\theta}}{\delta u})-\frac{\text{d}\mathbf{u}}{\text{d}t}||^{2}$
5:  $\theta\leftarrow\text{Update}(\theta,\nabla_{\theta}\mathcal{L})$
@@
Within this framework, the training is described in Algorithm 1. For Hamiltonian systems, we assume knowledge of the analytical form of $\mathcal{H}$ and $\delta\mathcal{H}/\delta u$, and labels for $\delta\mathcal{H}/\delta u$ or $\text{d}\mathbf{u}/\text{d}t$ can be calculated from training data using finite difference schemes. The training loss can either be evaluated in the functional form $\delta\mathcal{H}_{\theta}/\delta u$ or the temporal form $\mathcal{J}(\delta\mathcal{H}_{\theta}/\delta u)$; we implement the former since it more directly optimizes $\kappa_{\theta}$.
@@ 4.1 Toy Examples
To better understand and situate neural functionals, we examine its ability to model simple, analytically constructed functionals and compare its performance to other architectures. The experiment setup is as follows: given a functional $\mathcal{F}[u]$, random polynomials $u(x)=c_{0}x^{p}+c_{1}x^{p-1}+\ldots+c_{p-1}x+c_{p}$ are generated by uniformly sampling $\{c_{i}\in[a,b]:i=0,\ldots,p\}$. Therefore, the dataset consists of $N$ pairs of polynomials and evaluated functionals $(u^{n}(x),\mathcal{F}[u^{n}(x)])$ for $n=1,\ldots,N$. In practice, each function $u^{n}(x)$ is discretized at a set of points $\{x_{i}:i=1,\ldots,M\}$, such that it is represented as $\mathbf{u}^{n}=[u^{n}(x_{1}),\ldots,u^{n}(x_{L})]$, and $\mathcal{F}[\mathbf{u}^{n}]$ remains a scalar.
We consider two cases: a linear functional $\mathcal{F}_{l}[u]=\int_{x_{1}}^{x_{M}}u(x)*x^{2}\text{d}x$ and a nonlinear functional $\mathcal{F}_{nl}[u]=\int_{x_{1}}^{x_{M}}(u(x))^{3}\text{d}x$. In both cases, constructed polynomials $u^{n}(x)$ can be substituted and the definite integral can be analytically solved to use as training labels. Additionally, the functional derivatives are $\frac{\delta\mathcal{F}_{l}}{\delta u}=x^{2}$ and $\frac{\delta\mathcal{F}_{nl}}{\delta u}=3u^{2}$, which are used to evaluate the gradient performance of models. We restrict the degree of $u^{n}(x)$ to $p=2$ and sample $c_{i}\in[-1,1]$; 100 training and 10 validation samples are generated for each case. Coordinates $x_{i}$ are uniformly spaced on the domain $[-1,1]$ with $M=100$ points.
Two baselines are considered: (1) a MLP that takes concatenated inputs $[\mathbf{x},\mathbf{u}]\in\mathbb{R}^{2M}$ and projects to an output $\mathcal{F}_{\theta}\in\mathbb{R}$ through hidden layers, and (2) a FNO with mean-pooling at the final layer to evaluate an operator baseline. The FNO baseline takes inputs with spatial and channel dimensions, or stacked inputs $[\mathbf{x},\mathbf{u}]\in\mathbb{R}^{M\times 2}$ and projects to an output field $\mathbf{z}\in\mathbb{R}^{M}$, which is then pooled to $\mathcal{F}_{\theta}\in\mathbb{R}$. For simplicity, the proposed neural functional (NF) is used with a local MLP kernel and is linear for $\mathcal{F}_{l}$ and nonlinear for $\mathcal{F}_{nl}$. In the nonlinear case, FiLM is not used and $[x_{i},u_{i}]\in\mathbb{R}^{2}$ is concatenated as input to $\kappa_{\theta}$. Lastly, models are trained with the loss $\mathcal{L}=\sum_{i=1}^{n}||\mathcal{F}[u^{n}(x)]-\mathcal{F}_{\theta}(\mathbf{u}^{n},\mathbf{x})||_{2}^{2}$; training is supervised with the functional to evaluate if models can implicitly learn accurate derivatives.
@@
Metric MLP FNO NF MLP FNO NF MLP FNO NF
$\mathcal{F}_{l}[u]$ 1.0e-5 5.4e-4 9.8e-16 0.043 0.097 2.7e-14 31.21 31.53 0.21
$\delta\mathcal{F}_{l}/\delta u$ 0.081 0.24 7.0e-4 0.14 0.29 7.0e-4 2.64 2.68 0.033
$\mathcal{F}_{nl}[u]$ 0.12 0.026 0.0023 6131 5864 2126 315.9 281.0 62.3
$\delta\mathcal{F}_{nl}/\delta u$ 1.99 1.84 0.089 1684 1659 998 74.6 69.9 29.5
Table 1: Validation MSE is reported across different experiments with linear and nonlinear functionals. Errors are calculated in the scalar $\mathcal{F}[u]$ and gradient domains $\delta\mathcal{F}/\delta u$ to evaluate performance in fitting functionals and implicitly learning functional derivatives. Additionally, generalization to OOD inputs and unseen discretizations are evaluated. Parameter counts are: MLP (7.5K), FNO (10.9K) and NF (4.3K); each model takes roughly 30s to train until convergence.
To further evaluate generalization, after training we consider two cases: (1) OOD, where validation inputs $u(x)$ are sampled with polynomial coefficients $c_{i}\in[1,3]$ to model out-of-distribution functions, (2) Disc., where validation inputs $u(x)$ are uniformly discretized on a larger domain $x_{i}\in[-2,2]$ with a larger $\Delta x$, while keeping $M=100$. In each experiment, validation errors are reported with the metric $\frac{1}{n}\sum_{i=1}^{n}||\mathcal{F}[u^{n}]-\mathcal{F}_{\theta}(\mathbf{u}^{n},\mathbf{x})||_{2}^{2}$ or $\frac{1}{n}\sum_{i=1}^{n}||\frac{\delta\mathcal{F}}{\delta u^{n}}-\text{autograd}(F_{\theta},\mathbf{u}^{n})||_{2}^{2}$ to measure performance in fitting functionals and learning accurate functional derivatives.
Results are given in Table 1. We observe that neural functionals can accurately represent functionals and their derivatives, outperforming other architectures. Despite only training on scalars $\mathcal{F}[u]$, neural functionals can implicitly learn the correct functional derivatives, as plotted in Figure 2. This is new; even on toy problems, traditional architectures do not have well-behaved functional derivatives, as such, learning in the gradient domain for Hamiltonian systems would not be possible. Although the gradient performance is poor, conventional architectures can still fit functionals and the MSE for validation samples within the training distribution (Base) is low. When considering generalization performance, neural functionals can readily generalize to unseen input functions as well as unseen discretizations in the linear case. This empirically observed generalization is theoretically motivated; the Riesz representation theorem guarantees that a single kernel $\kappa_{\theta}$ can represent a linear functional $\mathcal{F}_{l}$ for all possible input functions $u(x)$ (e.g., polynomial, exponential, etc.). Generalization performance with nonlinear functionals is more challenging, as larger polynomial coefficients $c_{i}$ and larger coordinates $x_{i}$ can quickly increase $F_{nl}[u(x)]$ due to large exponents. Despite this, neural functionals still outperform other architectures in these domains, and learn accurate nonlinear functionals and their derivatives for samples within distribution.
@@
To evaluate the Hamiltonian Neural Functional (HNF) on PDE problems, we consider the 1D Advection and the 1D Korteweg–De Vries (KdV) equation. The Hamiltonians for each PDE are given as [miura_kdv, kdv_1]:
@@
which recovers the correct PDE for 1D Advection. A similar calculation can be done to recover the PDE for the 1D KdV equation ($u_{t}+uu_{x}+u_{xxx}=0$) from Hamilton’s equations by using $\frac{\delta\mathcal{H}_{kdv}}{\delta u}=-\frac{1}{2}u^{2}-u_{xx}$. Therefore, given a training dataset of spatiotemporal PDE data, the labels $\frac{\delta\mathcal{H}}{\delta u}$ can be computed using finite differences at every timestep $t$. The gradients of the HNF network $\nabla_{\mathbf{u}}(\mathcal{H}_{\theta}(\mathbf{x},\mathbf{u}^{t}))$ are then fitted to these labels.
@@
Data for the 1D Advection and KdV equations are generated from a numerical solver using random initial conditions [zhou2024maskedautoencoderspdelearners, brandstetter2022liepointsymmetrydata]:

### v2

# Hamiltonian Neural PDE Solvers through Functional Approximation

@@ Abstract
Designing neural networks within a Hamiltonian framework offers a principled way to ensure that conservation laws are respected in physical systems. While promising, these capabilities have been largely limited to discrete, analytically solvable systems. In contrast, many physical phenomena are governed by PDEs, which govern infinite-dimensional fields through Hamiltonian functionals and their functional derivatives. Building on prior work, we represent the Hamiltonian functional as a kernel integral parameterized by a neural field, enabling learnable function-to-scalar mappings and the use of automatic differentiation to calculate functional derivatives. This allows for an extension of Hamiltonian mechanics to neural PDE solvers by predicting a functional and learning in the gradient domain. We show that the resulting Hamiltonian Neural Solver (HNS) can be an effective surrogate model through improved stability and conserving energy-like quantities across 1D and 2D PDEs. This ability to respect conservation laws also allows HNS models to better generalize to longer time horizons or unseen initial conditions.
@@ 1 Introduction
Despite the long history of Hamiltonian mechanics, most current numerical and neural PDE solvers operate within a Newtonian framework. This has been very successful; forces and masses are easy to intuitively understand, and both numerical and neural solvers have found important, practical uses. However, motivated by natural conservation laws and symmetries that arise from Hamiltonian mechanics, we seek to investigate neural PDE solvers that operate in this framework. We find that the Hamiltonian framework brings inductive biases that improve model performance, especially when conserving energy-like quantities.
Extending Hamiltonian mechanics to neural PDE solvers requires both theoretical and implementation insights. In particular, most PDEs are used to describe continuum systems (fluids, waves, elastic bodies, etc.); as such, infinite-dimensional Hamiltonian mechanics are needed. In this setup, the Hamiltonian is defined as a functional that maps from input fields to a scalar, often interpreted as the energy of the system. The variational or functional derivative of the Hamiltonian then contains dynamical information that evolves the system. This introduces interesting modeling challenges; in particular, the need to approximate infinite-dimensional to scalar mappings and for the approximation to have functional derivatives.
In this paper, we consider models based on a learnable kernel integral, a framework that can provably approximate linear functionals and implicitly learn functional derivatives. Through automatic differentiation, functional derivatives can be obtained and optimized in a Hamiltonian framework to predict future PDE states. This allows the development of Hamiltonian Neural Solvers (HNS), which we evaluate on three PDE systems and find both high accuracy and generalization capabilities, which we hypothesize arise from inductive biases to conserve energy. Code and datasets for this work are released at https://github.com/anthonyzhou-1/hamiltonian_pdes.
@@ 2 Background
Many PDEs describe continuum systems, such as waves, fluids, and elastic bodies, where the positions and momenta of discrete bodies are not well defined. Despite this change, continuum systems can still conserve energy and be viewed in a Hamiltonian framework. Firstly, finite-dimensional vectors $\mathbf{q},\mathbf{p}\in\mathbb{R}^{n}$ become generalized to functions $u(x,t)$ defined on continuously varying coordinates $x\in\Omega$ and at a time $t$. The function $u\in\mathcal{F}(\Omega)$ describes the state of the system by assigning a scalar or vector for each coordinate (such as velocity, pressure, etc.). Secondly, the Hamiltonian is generalized from a function $\{\mathcal{H}(\mathbf{q},\mathbf{p}):\mathbb{R}^{2n}\rightarrow\mathbb{R}\}$ to a functional $\{\mathcal{H}[u]:\mathcal{F}(\Omega)\rightarrow\mathbb{R}\}$ that returns a scalar given an input function. Lastly, the symplectic matrix $J$ generalizes to a linear operator $\mathcal{J}$, and the gradient $\nabla_{\mathbf{u}}$ becomes the variational or functional derivative $\frac{\delta\mathcal{H}}{\delta u}\in\mathcal{F}(\Omega)$. This leads to Hamilton’s equations for infinite-dimensional systems, where $\epsilon\in\mathcal{F}(\Omega)$ is an arbitrary test function [1]:
@@
From this background we can see why neural networks are a natural choice for modeling discrete Hamiltonian systems; in this setting, the Hamiltonian is a function and its gradient is naturally calculated with automatic differentiation. In the infinite-dimensional setting, learning the Hamiltonian functional from data becomes less obvious.
3 Functional Approximation
@@
The development of functional approximators [2, 3, 4, 5] can be seen a result of the Riesz representation theorem, a fundamental result in functional analysis [6]. We give some intuition here, but provide a full proof of linear functional approximation in Appendix C.1. We start by observing that a functional can be equivalent to the inner product over suitable functions:
@@
Immediately, we can see that the Hamiltonian $\mathcal{H}[u]$ can potentially be expressed as an inner product $\left\langle u,\kappa_{\theta}\right\rangle$ over the input function $u$ and a learnable function $\kappa_{\theta}$. This changes the problem of approximating a functional to the problem of approximating a function, which neural networks can readily achieve through universal approximation theorems [7, 8].
A common inner product for arbitrary functions $u,v$ in a Hilbert space $H$ involves an integral: $\left\langle u,v\right\rangle=\int_{\Omega}u(x)v(x)\text{d}x$. One can check that this satisfies the basic properties of inner products and is also valid for vector-valued functions $u,v$. Based on this formulation, functionals can be parameterized by an Integral Kernel Functional (IKF):
@@
In this form, we observe a connection between the integral kernel functional and the integral kernel operator that neural operators are built on [9]. Specifically, the integral kernel operator is given by:
@@
$\mathcal{H}_{\theta}[u]=\int_{\Omega}\kappa_{\theta}(x,u(x))u(x)dx\approx\sum_{i}\kappa_{\theta}(x_{i},u(x_{i}))u(x_{i})\mu_{i}\Delta x$ (6)
with quadrature weights $\mu_{i}$. Interestingly, the full Riemann sum can be computed since the sum is only calculated once; in neural operator literature, approximating the integral with a full Riemann sum is usually too costly since the sum is evaluated for every query point. This results in approaches that truncate the sum [10, 11, 12] or represent it in Fourier space through the convolution theorem [13]. In our implementation, we evaluate different quadratures but find that the trapezoidal rule is enough to accurately estimate the integral.
To parameterize the kernel, we take the perspective from the Riesz representation theorem. In the linear case, the kernel is simply a function $\kappa_{\theta}(x)\in H$ that maps input coordinates to values in the codomain. This is exactly what a neural field is; therefore, we can inherit the substantial prior work on neural fields [14] to effectively parameterize $\kappa_{\theta}$.
One useful concept is the conditional neural field. In particular, the nonlinear kernel $\kappa_{\theta}(x,u(x))$ can be seen as a conditional neural field whereby the kernel output changes based on the input function $u(x)$. This concept can also be further extended to distinguish between local and global conditioning [15]; in neural field literature, the field can be conditioned on local information $u_{i}=u(x_{i})$ or global information $\mathbf{u}=[u(x_{0}),\ldots,u(x_{n})]$. Conditioning mechanisms can also be implemented in various ways. In the simplest case, the conditional information $u_{i}$ or $\mathbf{u}$ is concatenated to the input coordinate $x$ [16, 17, 18], however, we adopt a approach based on Feature-wise Linear Modulation (FiLM) [19].
Beyond conditioning mechanisms, an important consideration is the architecture of the kernel itself. A powerful neural field architecture is the sinusoidal representation network or SIREN [20]. SIRENs have shown good performance not only in fitting a neural field $\kappa_{\theta}$, but also in accurately representing gradients $\nabla\kappa_{\theta}$, which may help in fitting the functional derivative $\delta\mathcal{H}/\delta u$. With these architectural choices, a layer of the kernel $\kappa^{(l)}_{\theta}$ with local conditioning can be written as:
@@
Hamiltonian Neural Solvers
So far, integral kernel functionals are largely agnostic to the downstream task, only requiring that the input function $u$ be discretized at a set of points $(x_{i},u_{i})$. To use functional approximators as a PDE solver, we introduce the Hamiltonian Neural Solver (HNS). The time evolution of PDEs can be described by Hamilton’s equations, which depend on the operator $\mathcal{J}$ and the functional derivative $\delta\mathcal{H}/\delta u$. In Hamiltonian PDE systems, $\mathcal{J}$ is analytically derived or can be looked up, is guaranteed to be linear, and is usually simple. A common example in 1D is $\mathcal{J}=\partial_{x}$. Therefore, we choose to approximate $\mathcal{J}$ with a finite difference scheme, although it can also be learned with a neural operator. Lastly, we find that $\delta\mathcal{H}/\delta u$ can be computed directly with automatic differentiation. This is empirically shown in Section 4.1, but one way to see this is by considering the linear IKF. The gradient $\nabla_{\mathbf{u}}\sum_{i=1}^{n}\kappa_{\theta}(x_{i})u_{i}\>\mu_{i}\Delta x=[\kappa_{\theta}(x_{1}),\ldots,\kappa_{\theta}(x_{n})]$, up to a constant factor, which is the discretization of the learned function $\kappa_{\theta}(x)$. In this case, the functional derivative is approximated by an arbitrary function, which is the desired behavior.
Algorithm 1 Training a HNS
@@
[…]
@@
Within this framework, the training is described in Algorithm 1. For Hamiltonian systems, we assume knowledge of the analytical form of $\mathcal{H}$ and $\delta\mathcal{H}/\delta u$, and labels for $\delta\mathcal{H}/\delta u$ or $\text{d}\mathbf{u}/\text{d}t$ can be calculated from training data using numerical schemes. The training loss can either be evaluated in the functional form $\delta\mathcal{H}_{\theta}/\delta u$ or the temporal form $\mathcal{J}(\delta\mathcal{H}_{\theta}/\delta u)$; we implement the former since it more directly optimizes $\kappa_{\theta}$.
@@ 4.1 Toy Examples
To better understand and situate integral kernel functionals, we examine their ability to model simple, analytically constructed functionals and compare their performance to other architectures. The experimental setup is as follows: given a functional $\mathcal{F}[u]$, random polynomials $u(x)=c_{0}x^{p}+c_{1}x^{p-1}+\ldots+c_{p-1}x+c_{p}$ are generated by uniformly sampling $\{c_{i}\in[a,b]:i=0,\ldots,p\}$. Therefore, the dataset consists of $N$ pairs of polynomials and evaluated functionals $(u^{n}(x),\mathcal{F}[u^{n}(x)])$ for $n=1,\ldots,N$. In practice, each function $u^{n}(x)$ is discretized at a set of points $\{x_{i}:i=1,\ldots,M\}$, such that it is represented as $\mathbf{u}^{n}=[u^{n}(x_{1}),\ldots,u^{n}(x_{L})]$, and $\mathcal{F}[\mathbf{u}^{n}]$ remains a scalar.
We consider two cases: a linear functional $\mathcal{F}_{l}[u]=\int_{x_{1}}^{x_{M}}u(x)*x^{2}\text{d}x$ and a nonlinear functional $\mathcal{F}_{nl}[u]=\int_{x_{1}}^{x_{M}}(u(x))^{3}\text{d}x$. In both cases, constructed polynomials $u^{n}(x)$ can be substituted and the definite integral can be analytically solved to use as training labels. Additionally, the functional derivatives are $\frac{\delta\mathcal{F}_{l}}{\delta u}=x^{2}$ and $\frac{\delta\mathcal{F}_{nl}}{\delta u}=3u^{2}$, which are used to evaluate the gradient performance of models. We restrict the degree of $u^{n}(x)$ to $p=2$ and sample $c_{i}\in[-1,1]$; 100 training and 10 validation samples are generated for each case. Coordinates $x_{i}$ are uniformly spaced on the domain [1,1][-1,1] with $M=100$ points.
[…]
@@
Metric MLP FNO IKF MLP FNO IKF MLP FNO IKF
$\mathcal{F}_{l}[u]$ 2.47e-5 2.76e-4 3.00e-16 0.046 0.268 2.13e-7 49.3 49.7 0.737
$\delta\mathcal{F}_{l}/\delta u$ 0.083 0.066 1.15e-3 0.114 0.122 1.15e-3 2.64 2.70 0.077
$\mathcal{F}_{nl}[u]$ 0.029 0.016 2.05e-3 6709 6493 2908 381 322 55.4
$\delta\mathcal{F}_{nl}/\delta u$ 1.33 2.10 0.045 1730 1723 1079 92.1 88.1 35.5
Table 1: Validation MSE is reported across different experiments with linear and nonlinear functionals. Errors are calculated in the scalar $\mathcal{F}[u]$ and gradient domains $\delta\mathcal{F}/\delta u$ to evaluate performance in fitting functionals and implicitly learning functional derivatives. Additionally, generalization to OOD inputs and unseen discretizations are evaluated. Parameter counts are: MLP (7.5K), FNO (10.9K) and IKF (4.3K); each experiment is repeated 6 times and errors are reported as the average across seeds.
[…]
Results are given in Table 1. We observe that architectures based on an integral kernel can accurately represent functionals and their derivatives, outperforming other architectures. Despite only training on scalars $\mathcal{F}[u]$, IKFs can implicitly learn the correct functional derivatives, as plotted in Figure 2. This is new; even on toy problems, traditional architectures do not have well-behaved functional derivatives, as such, learning in the gradient domain for Hamiltonian systems would not be possible. Although the gradient performance is poor, conventional architectures can still fit functionals and the MSE for validation samples within the training distribution (Base) is low. When considering generalization performance, IKFs can readily generalize to unseen input functions as well as unseen discretizations in the linear case. This empirically observed generalization is theoretically motivated; the Riesz representation theorem guarantees that a single kernel $\kappa_{\theta}$ can represent a linear functional $\mathcal{F}_{l}$ for all possible input functions $u(x)$ (e.g., polynomial, exponential, etc.). Generalization performance with nonlinear functionals is more challenging, as larger polynomial coefficients $c_{i}$ and larger coordinates $x_{i}$ can quickly increase $F_{nl}[u(x)]$ due to large exponents. Despite this, IKFs still outperform other architectures in these domains, and learn accurate nonlinear functionals and their derivatives for samples within distribution.
@@
To evaluate the Hamiltonian Neural Solver (HNS) on PDE problems, we consider the 1D Advection and the 1D Korteweg–De Vries (KdV) equation. The Hamiltonians for each PDE are given as [21, 22]:
@@
which recovers the correct PDE for 1D Advection. A similar calculation can be done to recover the PDE for the 1D KdV equation ($u_{t}+uu_{x}+u_{xxx}=0$) from Hamilton’s equations by using $\frac{\delta\mathcal{H}_{kdv}}{\delta u}=-\frac{1}{2}u^{2}-u_{xx}$. Therefore, given a training dataset of spatiotemporal PDE data, the labels $\frac{\delta\mathcal{H}}{\delta u}$ can be computed using finite differences at every timestep $t$. The gradients of the HNS network $\nabla_{\mathbf{u}}(\mathcal{H}_{\theta}(\mathbf{x},\mathbf{u}^{t}))$ are then fitted to these labels.
@@
Data for the 1D Advection and KdV equations are generated from a numerical solver using random initial conditions [23, 24]: